40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Tue Sep 20 03:12:14 2011, Anamaria, Update, SUS, Jenne's Scripts started 
    Reply  Tue Sep 20 04:12:26 2011, Jenne, Update, SUS, Jenne's Scripts started 
       Reply  Tue Sep 20 09:44:44 2011, Jenne, Update, SUS, Jenne's Scripts started 
          Reply  Tue Sep 20 14:53:13 2011, Jenne, Update, SUS, Jenne's Scripts started 
Message ID: 5479     Entry time: Tue Sep 20 14:53:13 2011     In reply to: 5477
Author: Jenne 
Type: Update 
Category: SUS 
Subject: Jenne's Scripts started 

Quote:

Quote:

Quote:

I followed Jenne's instructions, ran the matrix filler script and then set the optics to freeswing. Someone has to burt resture and damp them in the morning.

 Thanks!  I'll give them a little more time, then restore things.

 I began restoring the optics at ~9:30am, so I have a full 6 hours of data, in case I need that much to separate the Pos/Side modes on some of the optics.  They are all damping again with their original matricies.

 So, clearly this was a kind of dumb idea.  There is nothing mechanical going on between our sensor inputs and our Pit/Pos/Yaw/Side DoF filter banks.  It's just math.  On the other hand, we now have a 3rd set of in-vac free swinging data, so I can (after all the suspensions are working) have a look at the drift in matrix elements over time.

In other news, after some meditation, and fitzing with DoF gain values, all of the IFO optics except for SRM now have their new input matricies, and are damping pretty nicely.  I need to go through and do an "eyeball" check to make sure that everything has a Q of ~5ish.  So far, I've kicked the optics, and watched that they damped fairly quickly, but I don't have a guesstimate of the Q's for each optic, for each DoF.

So, still to do:

Use another set of data and invert the SRM matrix DONE

Plug in the MC matricies, make sure they're okay. DONE

Check the Q's for all optics, all DoFs. 

ELOG V3.1.3-